# Dependence of the Mn sticking coefficient on Ga-rich, N-rich, and Ga/N-flux-free conditions in GaN grown by plasma-assisted molecular beam epitaxy

### v1

@@ III Results and Discussion
The areal densities of Mn in each Mn-doped region were determined by integrating the corresponding volume density profiles over depth. The integration limits were selected such that the Mn concentration reached the background signal. Based on the Mn shutter opening times (9 min for the N-rich and the Ga-rich GaN:Mn, and 34 s for the no-flux Mn doping) for each Mn-doped region, the resulting incorporated Mn fluxes were determined to be atoms cm-2 s-1, atoms cm-2 s-1, and atoms cm-2 s-1 for the N-rich, no-flux, and the Ga-rich conditions, respectively. These values yield relative Mn sticking coefficients, normalized to that under the N-rich condition, of 0.31 for the no-flux condition and 0.01 for the Ga-rich condition. The observed dependence of the Mn sticking coefficient on the Ga/N flux conditions can be understood by considering that Mn occupies cation (Ga) sites in the GaN crystal.Haider et al. (2003) Under N-rich growth conditions, Mn can readily access available cation sites. In contrast, under Ga-rich conditions, Mn competes with Ga for incorporation at cation sites. This difference leads to different Mn desorption rates and sticking coefficients. A Ga-adlayer-free GaN surface, achieved in the absence of both Ga and N fluxes during the Mn -doping, would result in a sticking coefficient that lies between the values observed under Ga-rich and N-rich conditions.
@@ V Acknowledgements
YongJin Cho thanks Oliver Bierwagen (PDI) for reviewing the manuscript.

### v2

@@ III Results and Discussion
The areal densities of Mn in each Mn-doped region were determined by integrating the corresponding volume density profiles over depth. The integration limits were selected such that the Mn concentration reached the background signal. Based on the Mn shutter opening times (9 min for the N-rich and the Ga-rich GaN:Mn, and 34 s for the no-flux Mn doping) for each Mn-doped region, the resulting incorporated Mn fluxes were determined to be atoms cm-2 s-1, atoms cm-2 s-1, and atoms cm-2 s-1 for the N-rich, no-flux, and the Ga-rich conditions, respectively. These values yield relative Mn sticking coefficients, normalized to that under the N-rich condition, of 0.31 for the no-flux condition and 0.01 for the Ga-rich condition. A similarly significant difference in Mn incorporation between Ga-rich and N-rich conditions has been reported previously.Kuroda et al. (2003) The observed dependence of the Mn sticking coefficient on the Ga/N flux conditions can be understood by considering that Mn occupies cation (Ga) sites in the GaN crystal.Haider et al. (2003) Under N-rich growth conditions, Mn can readily access available cation sites. In contrast, under Ga-rich conditions, Mn competes with Ga for incorporation at cation sites. This difference leads to different Mn desorption rates and sticking coefficients. A Ga-adlayer-free GaN surface, achieved in the absence of both Ga and N fluxes during the Mn -doping, would result in a sticking coefficient that lies between the values observed under Ga-rich and N-rich conditions.
@@ V Acknowledgements
YongJin Cho thanks Oliver Bierwagen (PDI) for reviewing the manuscript. This work was supported by the NSF FuSe project (Award 2329063), and by SUPREME, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA.